\newpage 
    \begin{thebibliography}{12}
      
\bibitem{Cun2015}
  Y. LeCun, Y. Bengio, and G. Hinton. "Deep learning." nature, pages 436--444, 2015.

\bibitem{Krizhevsky2012}
  A.~Krizhevsky, I.~Sutskever, and G.~E. Hinton, \newblock ``Imagenet classification with deep convolutional neural networks,'' \newblock in {NIPS}, vol.~60, pages 1097--1105, 2012.

  \bibitem{Jiu2015}
M.~Jiu and H.~Sahbi, ``Semi supervised deep kernel design for image annotation,'' in \emph{ICASSP}, 2015.



\bibitem{SzegedyCVPR2015}
C.~Szegedy et al. ``Going deeper with convolutions,'' in CVPR, 2015.
\bibitem{He2016}
K.~He, X.~Zhang, S.~Ren, and J.~Sun,
\newblock ``Deep residual learning for image recognition,''
\newblock in {CVPR}, 2016.

\bibitem{He2017}
He et al. "Mask r-cnn." Proceedings of ICCV, 2017.
\bibitem{HuangCVPR2017}
G.~Huang et al., ``Densely connected convolutional networks,''
\newblock in {CVPR}, 2017.
\bibitem{Jiu2016a}
M.~Jiu and H.~Sahbi, ``Laplacian deep kernel learning for image annotation,'' in   \emph{ICASSP}, 2016.

\bibitem{ross2015}
Girshick, Ross. "Fast r-cnn." Proceedings of the IEEE ICCV, 2015.

\bibitem{javad2017}
M. Javad, et al. "Fast YOLO: A fast you only look once system for real-time embedded object detection in video." arXiv:1709.05943 (2017).

\bibitem{ref00002}  H. Sahbi and F. Fleuret. Kernel methods and scale invariance using the triangular kernel. Diss. INRIA, 2004.

\bibitem{zaremba2014}
W. Zaremba, I. Sutskever, and O. Vinyals. "Recurrent neural network regularization." arXiv preprint arXiv:1409.2329 (2014).

\bibitem{zhang2020}
Z. Zhang, P. Cui, and W. Zhu. "Deep learning on graphs: A survey." In IEEE TKDE, 2020.
\bibitem{Jiu2017}
  M.~Jiu and H.~Sahbi, ``Nonlinear deep kernel learning for image annotation,'' \emph{IEEE  Transactions on Image Processing}, vol. 26(4), 2017.



\bibitem{indola2016}
Iandola, Forrest N., et al. "SqueezeNet: AlexNet-level accuracy with 50x fewer parameters and 0.5 MB model size." arXiv:1602.07360 (2016).

\bibitem{Chen2017}
  L-C. Chen et al. "Deeplab: Semantic image segmentation with deep convolutional nets, atrous convolution, and fully connected crfs." TPAMI 40.4 (2017): 834-848.
\bibitem{ref00001}  H. Sahbi and N. Boujemaa. "From coarse to fine skin and face detection." Proceedings of the eighth ACM international conference on Multimedia. 2000.


\bibitem{Long2015}
J. Long, E. Shelhamer and T. Darrell. "Fully convolutional networks for semantic segmentation." In IEEE CVPR, 2015.

\bibitem{indola2014}
Iandola, Forrest, et al. "Densenet: Implementing efficient convnet descriptor pyramids." arXiv:1404.1869 (2014).

\bibitem{Bruna2013}
J. Bruna, W. Zaremba, A. Szlam, Y. LeCun. Spectral networks and locally connected networks on graphs. arXiv preprint arXiv:1312.6203, 2013.
\bibitem{sahbiiccv17}
H. Sahbi. Coarse-to-fine deep kernel networks. Proceedings of the IEEE International Conference on Computer Vision, 1131-1139, 2017. 

\bibitem{Defferrard2016}
M. Defferrard, X. Bresson, P. Vandergheynst. Convolutional neural
networks on graphs with fast localized spectral filtering. In NIPS, pages 3844--3852, 2016.

\bibitem{ref000013} H. Sahbi, P. Etyngier, J-Y. Audibert, R. Keriven (2008, June). Manifold learning using robust graph laplacian for interactive image search. In 2008 IEEE Conference on Computer Vision and Pattern Recognition (pp. 1-8). IEEE.

\bibitem{Huang2018}
W. Huang, T. Zhang, Y. Rong, J. Huang. Adaptive sampling towards fast graph representation learning. In NIPS. pp. 4558-4567 (2018)


\bibitem{Kipf2016}
T.N. Kipf,  M. Welling. Semi-supervised classification with graph convolutional networks. arXiv preprint arXiv:1609.02907, 2016.

\bibitem{ref00003}   A. Dutta and H. Sahbi. "High order stochastic graphlet embedding for graph-based pattern recognition." arXiv preprint arXiv:1702.00156 (2017).

\bibitem{Henaff2015}
M. Henaff, J. Bruna, Y. LeCun. Deep convolutional networks on graphstructured
data. arXiv preprint arXiv:1506.05163 (2015)

\bibitem{Levie2018}
R. Levie, F. Monti, X. Bresson, M.M. Bronstein. Cayleynets: Graph
convolutional neural networks with complex rational spectral filters. IEEE
Transactions on Signal Processing 67(1), 97--109 (2018)

\bibitem{Li2018}
R. Li, S. Wang, F. Zhu, J. Huang. Adaptive graph convolutional neural
networks. In AAAI, 2018.
\bibitem{sahbipr2012}
F. Yuan, G-S. Xia, H. Sahbi, V. Prinet.  Mid-level Features and Spatio-Temporal Context for Activity Recognition. Pattern Recognition.  volume 45, number 12,  4182-4191, 2012
\bibitem{Zhuang2018}
Z. Chenyi and Q. Ma. Dual graph convolutional networks
for graph-based semi-supervised classification. Proceedings of WWW, 2018.

 \bibitem{ref00007} H. Sahbi, L. Ballan, G. Serra, A. DelBimbo (2012). Context-dependent logo matching and recognition. IEEE Transactions on Image Processing, 22(3), 1018-1031.


\bibitem{Chen2018}
J. Chen, T. Ma, C. Xiao. Fastgcn: fast learning with graph convolutional
networks via importance sampling. arXiv preprint arXiv:1801.10247 (2018)
\bibitem{sahbicbmi08}
H. Sahbi, JY. Audibert, J. Rabarisoa, R. Keriven. Object recognition and retrieval by context dependent similarity kernels. International Workshop on Content-Based Multimedia Indexing, 216-223, 2008.  

\bibitem{Gori2005}
M. Gori, G. Monfardini, F. Scarselli. A new model for learning in
graph domains. In IEEE IJCNN, vol. 2, pp. 729--734, 2005.
\bibitem{sahbiicip09}
M. Ferecatu, H. Sahbi. Multi-view object matching and tracking using canonical correlation analysis. 16th IEEE International Conference on Image Processing (ICIP), 2109-2112, 2009. 



\bibitem{Micheli2009}
A. Micheli. Neural network for graphs: A contextual constructive
approach. IEEE TNN 20(3), 498?511 (2009)

\bibitem{Scarselli2008}
F. Scarselli, M. Gori, A.C. Tsoi, M. Hagenbuchner, G. Monfardini. The graph neural network model. IEEE TNN 20(1), 61--80, 2008.

\bibitem{Wu2019}
Z. Wu, S. Pan, F. Chen, G. Long, C. Zhang, P.S. Yu. A comprehensive
survey on graph neural networks. arXiv:1901.00596 (2019).

\bibitem{ref00004}   H. Sahbi and X. Li. "Context-based support vector machines for interconnected image annotation." Asian Conference on Computer Vision. Springer, Berlin, Heidelberg, 2010.


\bibitem{Hamilton2017}
W. Hamilton, Z. Ying, J. Leskovec. Inductive representation learning
on large graphs. In NIPS. pp. 1024--1034 (2017)


\bibitem{WangICLR2018}
T. Wang, R. Liao, J. Ba, and S. Fidler. Nervenet: Learning structured policy with graph neural networks. ICLR, 2018

\bibitem{Loukas2020}
A. Loukas. What graph neural networks cannot learn: depth vs width. ICLR, 2020

\bibitem{Atamna}
A. Atamna, N. Sokolovska and J-C. Crivello. "A Principled Approach to Analyze Expressiveness and Accuracy of Graph Neural Networks." In ISIDA. Springer, 2020. 
\bibitem{sahbiicassp13a}
E. Benhaim, H. Sahbi, and G.  Vitte. "Designing relevant features for visual speech recognition." 2013 IEEE International Conference on Acoustics, Speech and Signal Processing. IEEE, 2013. 

\bibitem{lingsahbi2013}
L. Wang, H. Sahbi. Directed Acyclic Graph Kernels for Action Recognition. Proceedings of the IEEE International Conference on Computer Vision. 2013.

\bibitem{Sandeep2019}
 S. Kumar et al. "A unified framework for structured graph learning via spectral constraints." JMLR 21.22 (2020): 1-60.

\bibitem{Hanjun2017}
Khalil, Elias, et al. "Learning combinatorial optimization algorithms over graphs." In NIPS, 2017.

\bibitem{Marcelo2018}
Prates, Marcelo, et al. "Learning to solve NP-complete problems: A graph neural network for decision TSP." In  AAAI, vol. 33, 2019.

\bibitem{ref00005} N. Boujemaa, F. Fleuret, V. Gouet, H. Sahbi (2004, January). Visual content extraction for automatic semantic annotation of video news. In the proceedings of the SPIE Conference, San Jose, CA (Vol. 6).

\bibitem{Belkin2003}
M. Belkin and P. Niyogi. Lapl eigenmaps for dimensionality reduction and data representation. Neural computation 15.6 (2003): 1373-1396.

\bibitem{dong18}
X. Dong, D. Thanou, M. Rabbat, and P. Frossard, Learning graphs from data: A signal representation perspective, arXiv preprint arXiv:1806.00848, 2018.

\bibitem{Daitch}
S.I. Daitch, J.A. Kelner, and D.A. Spielman, Fitting a graph to vector data, in Proc. of ICML, 2009, pp. 201-208.
\bibitem{JiuPR2019}
M.~Jiu and H.~Sahbi, ``Deep representation design from deep kernel networks,''
  \emph{Pattern Recognition}, vol.~88, pp. 447--457, 2019.


\bibitem{Sardellitti16}
 S. Sardellitti, S. Barbarossa, and P. Di Lorenzo, Graph topology inference based on transform learning, in Proc. of the Global Conf. on Signal and Information Processing, 2016, pp. 356-360.

\bibitem{Sardellitti19}
S. Sardellitti et al., Graph topology inference based on sparsifying
transform learning. IEEE TSP, 67(7), 2019.
\bibitem{icassp2017b}
M. Jiu and H. Sahbi. Deep kernel map networks for image annotation. IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), 2016.


\bibitem{Valsesia18}
D. Valsesia, G. Fracastoro, and E. Magli, Sampling of graph signals via randomized local aggregations, arXiv preprint arXiv:1804.06182, 2018.

\bibitem{Kalofolias}
V. Kalofolias, How to learn a graph from smooth signals, in Proc. of the conf. on Artificial Intelligence and Statistics, 2016, pp. 920-929.

\bibitem{Egilmez}
H.E. Egilmez, E. Pavez, and A. Ortega, Graph learning from data under structural and laplacian constraints, arXiv preprint arXiv:1611.05181,2016.
\bibitem{lingsahbieccv2014}
 L. Wang, H. Sahbi. Nonlinear Cross-View Sample Enrichment for Action Recognition. European Conference on Computer Vision. Springer, 2014.


\bibitem{Chepuri}
S.P. Chepuri, S. Liu, G. Leus, and A.O. Hero, Learning sparse graphs under smoothness prior, in ICASSP, 2017.

\bibitem{Dong2016}
X. Dong, D. Thanou, P. Frossard, and P. Vandergheynst, Learning Laplacian matrix in smooth graph signal representations, IEEE TSP, 64(23), pages 6160--6173, 2016.

\bibitem{Pasdeloup2017}
B. Pasdeloup, V. Gripon, G. Mercier, D. Pastor, and M. G. Rabbat. Characterization and inference of graph diffusion processes from observations of stationary signals. IEEE TSIPN, 2017.
\bibitem{lingsahbiicip2014}
L. Wang, H. Sahbi. Bags-of-Daglets for Action Recognition. IEEE International Conference on Image Processing (ICIP), 2014.




\bibitem{Thanou2017}
D. Thanou, X. Dong, D. Kressner, and P. Frossard. Learning heat diffusion graphs. IEEE TSIPN, 3(3):484--499, 2017.

\bibitem{LeBars19}
  B. Le Bars et al.  Learning Laplacian Matrix from Bandlimited Graph Signals.  In ICASSP, 2019

\bibitem{kipf17} 
TN. Kipf, M. Welling. Semi-supervised classification with graph convolutional networks. In ICLR, 2017

\bibitem{ref00006}  H. Sahbi and F. Fleuret. Scale-invariance of support vector machines based on the triangular kernel. Diss. INRIA, 2002.


\bibitem{Yaguang19} 
Yaguang Li, Chuizheng Meng, Cyrus Shahabi, Yan Liu. Structure-informed Graph Auto-encoder for Relational Inference and Simulation. ICML, 2019

\bibitem{Kipf18} 
T Kipf, E Fetaya, KC Wang, M Welling, R Zemel. Neural Relational Inference for Interacting Systems. ICML, 2018

\bibitem{Sahbi2015}
H.~Sahbi, ``Imageclef annotation with explicit context-aware kernel maps,''   \emph{International Journal of Multimedia Information Retrieval}, pp.  113--128, 2015.



\bibitem{Ferran18} 
Ferran Alet, Adarsh K. Jeewajee, Maria Bauza, Alberto Rodriguez, Tomas Lozano-Perez, Leslie Pack Kaelbling. Graph Element Networks: adaptive, structured computation and memory. ICML, 2018

\bibitem{Alet18} 
Alet, F., Lozano-Perez, T., and Kaelbling, L. P. Modular meta-learning. In Proceedings of The 2nd Conference on Robot Learning, pp. 856--868, 2018.

\bibitem{Luca19} 
Luca Franceschi and Mathias Niepert and Massimiliano Pontil and Xiao He. Learning Discrete Structures for Graph Neural Networks. ICML, 2019. 

\bibitem{Sahbi2013icvs}
H.~Sahbi, ``Explicit context-aware kernel map learning for image annotation,''  in \emph{ICVS}, 2013.


\bibitem{Chen20} 
  Yu Chen et al., Deep Iterative and Adaptive Learning for Graph Neural Networks, AAAI DLGMA, 2020.

\bibitem{Li18} 
C. Li, Q. Zhong, D. Xie, and S. Pu. Co-occurrence feature learning from skeleton data for action recognition and detection with hierarchical aggregation. arXiv preprint arXiv:1804.06055, 2018.


\bibitem{ref00009}  H. Sahbi, D. Geman, N. Boujemaa. "Face detection using coarse-to-fine support vector classifiers." Proceedings. International Conference on Image Processing. Vol. 3. IEEE, 2002.

\bibitem{Chenyi18} 
  Z. Chenyi and Q. Ma. Dual graph convolutional networks for graph-based semi-supervised classification. Proceedings of WWW, 2018.

\bibitem{YLi2018}
Y. Li, R. Yu, C. Shahabi, and Y. Liu, Diffusion convolutional recurrent
neural network: Data-driven traffic forecasting. In Proc. of ICLR, 2018.

\bibitem{Yasunori2005}
N. Yasunori. A note on Riemannian optimization methods on the Stiefel and
the Grassmann manifolds. In NOLTA, vol 1, pages 349--352, 2005.

\bibitem{sahbikpca06}
H Sahbi.  Kernel PCA for similarity invariant shape recognition. Neurocomputing 70 (16-18), 3034-3045

\bibitem{HuangAAAI2017}
L. Huang, X. Liu, B. Lang, A. W. Yu, Y. Wang, and B. Li. Orthogonal weight normalization: Solution to optimization over multiple dependent Stiefel manifolds in deep neural networks. In AAAI, 2017.

\bibitem{Ankita2019}
A. Shukla, S. Bhagat, S. Uppal, S. Anand, P. Turaga. PrOSe: Product of Orthogonal Spheres Parameterization for Disentangled Representation Learning. In BMVC, 2019.  

\bibitem{thiemert2005applying}
S.~Thiemert, H.~Sahbi, and M.~Steinebach, ``Applying interest operators in  semi-fragile video watermarking,'' in \emph{Security, Steganography, and   Watermarking of Multimedia Contents VII}, vol. 5681.\hskip 1em plus 0.5em   minus 0.4em\relax International Society for Optics and Photonics, 2005, pp.  353--363.

\bibitem{Adam2014}
 D.P. Kingma, and J. Ba. "Adam: A method for stochastic optimization." arXiv preprint arXiv:1412.6980 (2014)

\bibitem{Sahbi2011}
H.~Sahbi, J.-Y. Audibert, and R.~Keriven, ``Context-dependent kernels for  object classification,'' \emph{IEEE Transactions on Pattern Analysis and   Machine Intelligence}, vol.~33, pp. 699--708, 2011.


\bibitem{garcia2018}
  G. Garcia-Hernando, S. Yuan, S. Baek, and T.-K. Kim. First Person Hand Action Benchmark with RGB-D Videos and 3D Hand Pose Annotations. In CVPR, 2018.

  \bibitem{refref10} 
  C. Feichtenhofer, A. P., and A. Zisserman. Convolutional Two-Stream Network Fusion for Video Action Recognition. In CVPR,  2016.

\bibitem{sahbijstars17}
Q. Oliveau, H. Sahbi. Learning attribute representations for remote sensing ship category classification.  IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing, 2017. 

  \bibitem{refref39}
  E.Ohn-Barand, M.M.Trivedi. Hand Gesture Recognition in Real Time for Automotive Interfaces: A Multimodal Vision- Based Approach and Evaluations. IEEE TITS, 15(6), pages 2368--2377, 2014.

\bibitem{ref00008}  N. Bourdis, D. Marraud, H. Sahbi. "Constrained optical flow for aerial image change detection." 2011 IEEE International Geoscience and Remote Sensing Symposium. IEEE, 2011.

  
   \bibitem{refref40} 
  O. Oreifej and Z. Liu. HON4D: Histogram of Oriented 4D Normals for Activity Recognition from Depth Sequences. In CVPR, pages 716--723, 2013.
   \bibitem{refref41}  
  H. Rahmani and A. Mian. 3D Action Recognition from Novel Viewpoints. In CVPR, pages 1506--1515, 2016.


\bibitem{sahbijmlr06}
H. Sahbi, D. Geman. A hierarchy of support vector machines for pattern detection. Journal of Machine Learning Research 7.Oct (2006): 2087-2123.



  \bibitem{Zhua2016} 
  W. Zhu et al., Co-occurrence feature learning for skeleton based action recognition using regularized deep LSTM networks In AAAI, 2016.

  \bibitem{refref59} 
  M. Zanfir et al., The Moving Pose: An Efficient 3D Kinematics Descriptor for Low-Latency Action Recognition and Detection. In ICCV, 2013.

  \bibitem{Vemulapalli2014}
  R. Vemulapalli, F. Arrate, and R. Chellappa. Human action recognition by representing 3D skeletons as points in a Lie group. In IEEE CVPR, 2014.

  \bibitem{ref000010}  H. Sahbi. "CNRS-TELECOM ParisTech at ImageCLEF 2013 Scalable Concept Image Annotation Task: Winning Annotations with Context Dependent SVMs." CLEF (Working Notes). 2013.

  \bibitem{Du2015}  
  Y. Du et al., Hierarchical recurrent neural network for skeleton based action recognition. In IEEE CVPR, 2015.


\bibitem{sahbiigarss12a}
N. Bourdis, D. Marraud, H. Sahbi.  "Camera pose estimation using visual servoing for aerial video change detection." 2012 IEEE International Geoscience and Remote Sensing Symposium. IEEE, 2012.


  \bibitem{refref61}  
  X. Zhang et al., Efficient Temporal Sequence Comparison and Classification Using Gram Matrix Embeddings on a Riemannian Manifold. In CVPR, 2016.

  \bibitem{ref000011}  H. Sahbi. "A particular Gaussian mixture model for clustering and its application to image retrieval." Soft Computing 12.7 (2008): 667-676.

 \bibitem{refref11}  
  G. Garcia-Hernando and T.-K. Kim. Transition Forests: Learning Discriminative Temporal Transitions for Action Recognition. In CVPR, 2017.


\bibitem{ref000012} H. Sahbi, J-Y. Audibert, R. Keriven. "Graph-cut transducers for relevance feedback in content based image retrieval." 2007 IEEE 11th International Conference on Computer Vision. IEEE, 2007.

  
  \bibitem{refref18}   
  J. Hu et al., Jointly Learning Heterogeneous Features for RGB-D Activity Recognition. In CVPR, 2015

  \bibitem{Huangcc2017} 
  Z. Huang and L. V. Gool. A Riemannian Network for SPD Matrix Learning. In AAAI, 2017


\bibitem{Vo2012}
P.~Vo and H.~Sahbi, ``Transductive kernel map learning and its application to image annotation,'' in \emph{BMVC}, 2007.

\bibitem{ref23}
 Z. Huang, J. Wu, and L. V. Gool. Building Deep Networks on Grassmann Manifolds. In AAAI, 2018

\bibitem{ref000014}   T. Napoléon and H. Sahbi. "From 2D silhouettes to 3D object retrieval: contributions and benchmarking." EURASIP Journal on Image and Video Processing 2010 (2010): 1-17.

  
\end{thebibliography}

\documentclass[12pt,journal,compsoc,onecolumn]{IEEEtran}

\usepackage{amsmath,graphicx,cleveref,caption,subcaption,dsfont,multirow,multicol,amssymb,amsfonts,mathtools,pifont}
\usepackage[noadjust]{cite}

\newtheorem{proposition}{Proposition}
\newtheorem{definition}{Definition}
\newtheorem{proof}{Proof}

\newcommand{\cmark}{\ding{51}}
\newcommand{\xmark}{\ding{55}}

\def\JJ{{\bf J}}

\def\U{{\mathbf U}} 
\def\V{{\mathbf V}} 
\def\A{{\mathbf \Lambda}}

\def\A{{\bf A}}

\def\I{{\bf I}}

\def\U{{\bf U}}
\def\W{{\bf W}}
\def\S{{\cal S}}
\def\N{{\cal N}}
\def\G{{\cal G}}
\def\V{{\cal V}}
\def\E{{\cal E}}
\def\F{{\cal F}}
\def \v{{\bf vec}} 
\def \J{{E}}
\def\tr{{\bf Tr}} 
\def\D{{\bf D}}
\def\M{{\bf M}}

\DeclarePairedDelimiter\floor{\lfloor}{\rfloor}

\title{Skeleton-based Hand-Gesture Recognition with Lightweight Graph Convolutional Networks}

\author{Hichem Sahbi \\ CNRS Sorbonne University}

\begin{document}
\maketitle
\begin{abstract}
  Graph convolutional networks (GCNs) aim at extending deep learning to arbitrary irregular domains, namely graphs. Their success is highly dependent on how the topology of input graphs is defined and most of the existing GCN architectures rely on predefined or handcrafted graph structures.\\ 
  In this paper, we introduce a novel method that learns the topology (or connectivity) of input graphs as a part of GCN design. The main contribution of our method resides in building an orthogonal connectivity basis that optimally aggregates nodes, through their neighborhood, prior to achieve convolution. Our method also considers a stochasticity criterion which acts as a regularizer that makes the learned basis and the underlying GCNs lightweight while still being highly effective. Experiments conducted on the challenging task of skeleton-based hand-gesture recognition show the high effectiveness of the learned GCNs w.r.t. the related work. \\
  
  {\noindent {\bf Keywords.} Graph convolutional networks, lightweight connectivity design, skeleton-based hand-gesture recognition.}
\end{abstract}

\section{Introduction}
\label{sec:intro}

Deep learning is currently witnessing a major interest in different fields including image processing and pattern recognition \cite{Cun2015}.  Its principle consists in learning multi-layered convolutional, pooling and fully connected operations that extract representations which capture low, mid and high-level characteristics of patterns while maximizing their classification performances. Most of the existing deep learning architectures \cite{Krizhevsky2012,SzegedyCVPR2015,He2016,HuangCVPR2017,He2017,ross2015,javad2017,zaremba2014,zhang2020,indola2016,Chen2017,Long2015,indola2014,Jiu2015,Jiu2016a,Jiu2017,JiuPR2019,icassp2017b,sahbiiccv17} are targeted to vectorial data; i.e., data sitting on top of regular domains including images.  However, other data require extending deep learning to irregular domains (namely graphs  \cite{zaremba2014,Bruna2013,Defferrard2016,Huang2018,Kipf2016,Sahbi2011,Sahbi2013icvs,Sahbi2015,sahbicbmi08
}) such as skeletons in action recognition. While convolutional operations on regular domains are well defined, their extension to irregular ones (i.e., graphs) is generally ill-posed and remains a major challenge. \\

\indent Two different categories of GCNs exist in the literature, spatial and spectral \cite{Gori2005,Micheli2009,Scarselli2008,Wu2019,Hamilton2017,zhang2020}. Spatial methods achieve node aggregations prior to apply convolutions using inner products while spectral techniques rely on the well defined graph Fourier transform \cite{Bruna2013,Defferrard2016,Henaff2015,Kipf2016,Levie2018,Li2018,Zhuang2018,Chen2018,Huang2018}. Whereas spatial methods are known to be effective compared to spectral ones, their success is highly dependent on the topology of input graphs, and most of the existing solutions rely on handcrafted or predefined graph structures using similarities or the inherent properties of the targeted applications \cite{WangICLR2018,Loukas2020,Atamna,sahbiicip09,lingsahbiicip2014,lingsahbieccv2014,sahbiicassp13a,lingsahbi2013,sahbipr2012}  (e.g.,  node relationships in social networks, edges in 3D modeling, etc). These structures are usually powerless to capture the most prominent relationships between nodes as their design is agnostic to the targeted application.  For instance, when considering node relationships in skeletons, these links capture the anthropometric characteristics of individuals which are useful for their identification, while other connections, yet to infer, are important for recognizing their actions. Hence, in spite of being relatively effective, the potential of these GCN methods is not fully explored as the setting of their graphs is either oblivious to the tasks at hand or achieved using the tedious cross validation. \\

\indent Graph inference  is generally ill-posed, NP-hard  \cite{Sandeep2019, Hanjun2017,Marcelo2018} and most of the existing approaches rely on constraints (similarity, smoothness, sparsity, band-limitedness, etc. \cite{Belkin2003,dong18,Daitch,Sardellitti16,LeBars19,Sardellitti19,Valsesia18,Kalofolias,Egilmez,Chepuri,Dong2016}) for its conditioning~\cite{dong18,LeBars19,Sardellitti19,Chepuri,Dong2016,Pasdeloup2017,Thanou2017}. Particularly in GCNs, recent advances aim at defining graph topology that best fits a given task \cite{Yaguang19,Kipf18,Ferran18,Alet18,Luca19,Chen20,Li18,Chenyi18}. For instance, \cite{Luca19} proposes a graph network for semi-supervised classification that learns graph topology with sparse structure given a cloud of points; node-to-node connections are modeled with a joint probability distribution on Bernoulli random variables whose parameters are found using bi-level optimization. A computationally more efficient variant is introduced in \cite{Chen20} using a weighted cosine similarity and edge thresholding. Other solutions make improvement w.r.t. the original GCNs \cite{Kipf2016} by exploiting symmetric matrices \cite{Li18} and discovering hidden structural relations (unspecified in the original graphs), using a so-called residual graph adjacency matrix and by learning a distance function over nodes. The work in \cite{Chenyi18} introduces a dual architecture with two parallel graph convolutional layers sharing the same parameters, and considers a normalized adjacency and a positive point-wise mutual information matrix to capture node co-occurrences through random walks sampled from graphs. \\

\indent In this paper, we introduce  a novel framework that designs graphs as a part of end-to-end GCN learning. Our design principle is based on the minimization of a constrained loss  whose solution corresponds not only to the convolutional parameters of GCNs but also the underlying  adjacency matrices that capture the topology of input graphs. Our contribution in this paper differs from the aforementioned  related work in multiple aspects; on the one hand, in contrast to many existing methods -- e.g., \cite{YLi2018} which consider a single adjacency matrix shared through power series -- the matrix operators designed in our contribution are non-parametrically learned and this provides more flexibility to our design. On the other hand, constraining these matrices, through orthogonality and stochasticity,  allows achieving structured regularization that mitigates overfitting and allows learning lightweight GCN architectures\footnote{\scriptsize Adjacency matrices learned,  in the related work, are usually fully dense and this introduces a lot of latency in the underlying GCNs, especially when handling large scale databases (even with reasonable size graphs).}; in contrast to non structured lightweight network design  (e.g., magnitude pruning), our proposed method  (i) captures (through orthogonality and stochasticity) the structural relationships between   parameters in the learned GCNs, and (ii) maintains completeness and minimality of the learned representations by finding the most discriminating and lightweight GCNs as also supported in our experiments.  

\section{Learning lightweight connectivity} 
 
Let $\S=\{\G_i=(\V_i, \E_i)\}_i$ denote a collection of graphs with $\V_i$, $\E_i$ being respectively the nodes and the edges of $\G_i$. Each graph $\G_i$ (denoted for short as $\G=(\V, \E)$) is endowed with a signal $\{\psi(u) \in \mathbb{R}^s: \ u \in \V\}$ and associated with an adjacency matrix $\A$ with each entry  $\A_{uu'}>0$ iff $(u,u') \in \E$ and $0$ otherwise. GCNs aim at learning a set of filters $\F=\{g_\theta=(\V_\theta,\E_\theta)\}_{\theta=1}^C$ that define convolution on $n$ nodes of $\G$ (with $n=|\V|$) as
 \begin{equation}\label{matrixform} 
(\G \star \F)_\V = f\big(\A \  \U^\top  \   \W\big), 
 \end{equation} 
 \noindent  here $^\top$ stands for transpose,  $\U \in \mathbb{R}^{s\times n}$  is the  graph signal, $\W \in \mathbb{R}^{s \times C}$  is the matrix of convolutional parameters corresponding to the $C$ filters and  $f(.)$ is a nonlinear activation applied entrywise. In Eq.~\ref{matrixform}, the input signal $\U$ is projected using $\A$ and this provides for each node $u$, the  aggregate set of its neighbors. When $\A$ is common to  all graphs\footnote{\scriptsize e.g., when considering a common graph structure for all actions in videos.}, entries of $\A$ could be handcrafted or learned so Eq.~(\ref{matrixform}) implements a convolutional block with two layers; the first one aggregates signals in $\N(\V)$ (sets of node neighbors) by multiplying $\U$ with $\A$ while the second layer achieves convolution by multiplying the resulting aggregates with the $C$ filters in $\W$.
 \subsection{Orthogonality-driven connectivity}\label{ortho}
Learning  multiple adjacency matrices (denoted as $\{\A_k\}_{k=1}^K$) allows us to capture different contexts and graph topologies when achieving aggregation and convolution.  With multiple matrices $\{\A_k\}_k$ (and associated convolutional filter parameters $\{\W_k\}_k$),  Eq.~\ref{matrixform} is updated as  

\begin{equation}\label{matrixform2} 
  (\G \star \F)_\V = f\bigg(\sum_{k=1}^K \A_k   \U^\top     \W_k\bigg).
 \end{equation} 
 If aggregation produces,  for a given $u \in \V$,  linearly dependent vectors ${\cal X}_u= \{\sum_{u'} \A_{kuu'}. \psi(u')\}_k$, then convolution will also generate  linearly dependent representations with an overestimated number of training  parameters in the null space of ${\cal X}_u$. Besides, the tensor $\{\A_k\}_k$ used for aggregation,  may also generate overlapping and redundant contexts.\\  Provided that  $\{\psi(u')\}_{u' \in \N_r(u)}$ are  linearly independent, the sufficient condition that makes vectors in ${\cal X}_u$ linearly independent reduces to  constraining $(\A_{kuu'})_{k,u'}$ to lie on the  Stiefel manifold  (see for instance \cite{Yasunori2005,HuangAAAI2017,Ankita2019}) defined as $V_K(\mathbb{R}^{n})=\{ \M \in  \mathbb{R}^{K \times n}: \M \, \M^\top =\I_K\}$  (with $\I_K$ being the $K \times K$ identity matrix) which thereby guarantees  orthonormality and   minimality of  $\{\A_1,\dots,\A_K\}$\footnote{\scriptsize Note that $K$ should not exceed the rank of $\big\{\psi(u')\big\}_{u' \in \N_r(u)}$ which is upper bounded by $\min(|{\cal V}|,s)$; $s$ is again the dimension of the graph signal.}. A less compelling condition is orthogonality, i.e.,  $\langle \A_k,\A_{k'} \rangle_F=0$ and $\A_{k}\geq {\bf 0}_{n\times n}$, $\A_{k'}\geq {\bf 0}_{n \times n}$,  $\forall k \neq k'$ --- with $\langle, \rangle_F$ being the Hilbert-Schmidt (or Frobenius) inner product defined as  $\langle \A_k,\A_{k'} \rangle_F=\tr(\A_k^\top\A_{k'})$ --- and this equates $\A_k\odot \A_{k'}= {\bf 0}_{n\times n} $,  $\forall k\neq k'$ with $\odot$ denoting the entrywise hadamard product and  ${\bf 0}_{n \times n}$ the $n \times n$ null matrix. \\ 

 \indent Considering orthogonality (as discussed above), the tensor  $\{\A_k\}_k$ and $\W=\{ \W_k \}_k$ are learned as 
\begin{equation}\label{matrixform3} 
  \begin{array}{lll}
\displaystyle {\displaystyle \min}_{\{\A_k \geq 0\}_k,\W} \ \ \  & \displaystyle   E\big(\A_1,\dots,\A_K;\W\big) & \\
& & \\
\displaystyle  {\textrm{s.t.}} &  \A_k\odot \A_{k}> {\bf 0}_{n \times n}  &  \\
                                                          &  \A_k\odot \A_{k'}= {\bf 0}_{n \times n}  &      \forall   k, k' \neq k \\
     &  {\bf 1}_n^\top \A_k =  {\bf 1}_n^\top.
\end{array}
\end{equation}
being $E$ the cross entropy loss and  ${\bf 1}_n^\top$ a vector of $n$ ones. In the above minimization problem, the first and the second constraints correspond to orthogonality  while the third one to column-stochasticity. The latter is added in order to ensure that all of the entries in $\A_k$ are positive and each column sums to one; i.e., each matrix $\A_k$ models a Markov chain whose $i$-th row and $j$-th column provides the probability of transition from one node $u_j$ to $u_i$ in $\G$. Note that orthogonality (as designed subsequently) allows learning sparse adjacency matrices while column-stochasticity provides extra sparsity and acts as a structured regularizer that enhances further the generalization power of the learned GCNs\footnote{\scriptsize Without stochasticity, one has to consider a normalization layer (with extra parameters), especially on graphs with heterogeneous degrees in order to reduce the covariate shift and distribute the transition probability evenly through nodes before achieving convolutions.}.
\subsection{Optimization} 

A natural approach to solve Eq.~(\ref{matrixform3}) is to iteratively and alternately  minimize over one matrix while keeping all the others fixed. However --- and besides the non-convexity of the loss --- the feasible set formed by these $O(K^2)$ bi-linear constraints is not convex w.r.t $\{\A_k\}_k$. Moreover, this iterative procedure is computationally expensive as it requires solving multiple instances of constrained projected gradient descent and the number of necessary iterations to reach convergence is large in practice. All these issues make solving this  problem  challenging and computationally intractable even for reasonable values of $K$ and $n$. In what follows, we investigate a  workaround that   optimizes these matrices while guaranteeing their  orthogonality  and stochasticity as a part  of optimization.\\

\noindent {\bf Orthogonality.} Let $\exp(\gamma \hat{\A}_{k}) \oslash (\sum_{r=1}^K \exp(\gamma \hat{\A}_{r}))$ be a softmax reparametrization of $\A_{k}$, with  $\oslash$ being the entrywise hadamard division and  $\{\hat{\A}_k\}_k$ free parameters in $\mathbb{R}^{n \times n}$, it becomes possible to implement orthogonality by choosing large values of  $\gamma$ to make this softmax {\it crisp}; i.e., only one entry $\A_{kij}\gg 0$ while all others $\{\A_{k'ij}\}_{k'\neq k}$ vanishing thereby leading to $\A_k\odot \A_{k'}= {\bf 0}_{n\times n}$,  $\forall   k, k' \neq k$. By plugging this {\it crispmax} reparametrization into Eq.~\ref{matrixform3}, the gradient of the loss $\J$ (now w.r.t   $\{\hat{\A}_k\}_k$)  is updated using the chain rule as
\begin{equation}\label{eq00001111}
  \begin{array}{lll}
\displaystyle      & \displaystyle \frac{\partial \J}{\partial \v(\{\hat{\A}_k\}_k)} &= \displaystyle  \JJ_\textrm{orth} .   \frac{\partial \J}{\partial \v({\{{\A}_k\}_k})},
 \end{array}
                                                                                                                                                                                           \end{equation} 
being $\v(\{\A_k\}_k)$ a vectorization of $\{\A_k\}_k$ and $({\bf i},{\bf j})=(kij,k'i'j')$ an entry of the Jacobian $\JJ_{\textrm{orth}}$ as  
\begin{equation}\label{eq00001112}
  \begin{array}{lll}
   &\displaystyle  \left\{ \begin{array}{ll} \gamma {\A}_{kij}.(1 - \A_{k ij}) &  {\footnotesize \textrm{if} \ k=k', i=i', j=j'} \\ -\gamma {\A}_{kij}.\A_{k' ij} &  {\small \textrm{if} \ k \neq k', i=i', j=j'}   \\ 0 &  \textrm{\small  otherwise,}  \end{array}\right. 
                                                                                                                                                                                         \end{array}
                                                                                                                                                                                           \end{equation} 
                                                                                                                                                                                           \noindent here $\frac{\partial \J}{\partial \v({\{{\A}_k\}_k})}$   is obtained from layerwise gradient backpropagation. However, with this reparametrization, large values of  $\gamma$  may lead to numerical instability when evaluating the exponential. We circumvent this by choosing $\gamma$ that satisfies $\epsilon$-orthogonality: a surrogate property defined subsequently. \\
                                                                                                                                                                                           
\begin{definition}[$\epsilon$-orthogonality]  A basis  $\{\A_k\}_k$ is $\epsilon$-orthogonal if $\A_k\odot \A_{k'} \leq \epsilon  \ \mathds{1}_{n\times n}$, $\forall k,k' \neq k$,  with $\mathds{1}_{n \times n}$ being the $n \times n$ unitary matrix.
\end{definition}

Considering the above definition, (nonzero) matrices belonging to an $\epsilon$-orthogonal basis are linearly independent w.r.t $\langle .,. \rangle_F$ (provided that $\gamma$ is sufficiently large) and  hence this basis  is also minimal. The following proposition provides a tight lower bound on $\gamma$ that satisfies $\epsilon$-orthogonality. \\

\begin{proposition} [$\epsilon$-orthogonality bound] Consider $\{\A_{kij}\}_{ij}$ as the entries of the crispmax reparametrized  matrix  $\A_k$ defined as  $\exp(\gamma \hat{\A}_{k}) \oslash \big(\sum_{r=1}^K \exp(\gamma \hat{\A}_{r})\big)$.  Provided that  $\exists \delta>0:$ $\forall  i,j,\ell'$, $\exists !\ell$,  $\hat{\A}_{\ell ij} \geq  \hat{\A}_{\ell' ij}+\delta$ (with $\ell' \neq \ell$) and if $\gamma$ is at least {$$\displaystyle \frac{1}{\delta} \ln\bigg(\frac{K \sqrt{(1-2\epsilon)}}{1-\sqrt{(1-2\epsilon)}}+1\bigg)$$}
then  $\{\A_1,\dots,\A_K\}$ is $\epsilon$-orthogonal.  
\end{proposition}
\begin{proof}  
For any entry $i,j$, one may find $\ell$, $\ell'$ in $\{1,\dots,K\}$ (with $\ell \neq \ell'$) s.t.  $(\A_k \odot \A_{k'})_{ij} $ 
{\hspace*{-1cm} \begin{equation*}
\begin{array}{lll}
\displaystyle  &\leq &  \displaystyle  (\A_\ell \odot \A_{\ell'})_{ij}  \\ 
& & \\
&= &   \frac{1}{2} (\A_{\ell ij}^2 + \A_{\ell' ij}^2) -   \frac{1}{2} (\A_{\ell ij} -\A_{\ell' ij})^2     \\ 
& & \\
\displaystyle    &\leq&     \frac{1}{2}-   \frac{1}{2} (\A_{\ell ij} -\A_{\ell' ij})^2     \\ 
\displaystyle   &= &  \frac{1}{2}-  \frac{1}{2} \bigg( \displaystyle\frac{\exp(\gamma \hat{\A}_{\ell ij})-\exp(\gamma \hat{\A}_{\ell' ij})}{\exp(\gamma \hat{\A}_{\ell ij})+\exp(\gamma \hat{\A}_{\ell' ij})+\sum_{r=3}^K \exp(\gamma \hat{\A}_{rij})}\bigg)^2  \\ 
& &\\
\displaystyle &\leq &   \frac{1}{2} - \frac{1}{2}\bigg(\displaystyle\frac{\exp(\gamma \hat{\A}_{\ell ij}) -\exp(\gamma \hat{\A}_{\ell' ij})}{\exp(\gamma \hat{\A}_{\ell ij}) +(K-1)\exp(\gamma \hat{\A}_{\ell' ij})}\bigg)^2  \\
& & \\
\displaystyle & \leq  &   \frac{1}{2}  -  \frac{1}{2} \bigg(\displaystyle \frac{1}{1+ \frac{K}{\exp(\gamma \delta )-1}}\bigg)^2.
 \end{array}
\end{equation*}
The sufficient condition is to choose $\gamma$ such as  
\begin{equation*}
 \frac{1}{2}  -  \frac{1}{2} \bigg[\displaystyle \frac{1}{1+ \frac{K}{\exp(\gamma \delta )-1}}\bigg]^2 \leq \epsilon \implies \displaystyle \gamma \geq\displaystyle \frac{1}{\delta} \ln\bigg(\frac{K \sqrt{(1-2\epsilon)}}{1-\sqrt{(1-2\epsilon)}}+1\bigg).
 \end{equation*} 
 }
\begin{flushright}
$\blacksquare$
\end{flushright}

\end{proof} 
Following the above  proposition, setting $\gamma$ to the above lower bound guarantees $\epsilon$-orthogonality; for instance, when $K=2$, $\delta=0.01$ and provided that  $\gamma \geq 530$, one may obtain  $0.01$-orthogonality which is almost a strict orthogonality. This property is satisfied as long as one slightly disrupts the entries of  $\{\hat{\A}_k\}_k$ with random noise during training\footnote{\scriptsize whatever  the range of entries in these matrices $\{\hat{\A}_k\}_k$.}. However, this may still lead to another limitation; precisely,  bad local minima are observed  due to an {\it early} convergence to crisp adjacency matrices. We prevent this by steadily annealing the temperature $1/\gamma$  of the softmax through training epochs (using $\frac{\gamma.\textrm{epoch}}{\textrm{max\_epochs}}$ instead of $\gamma$) in order to make optimization focusing first on the loss, and then as optimization evolves, temperature cools down and allows reaching the aforementioned lower bound (thereby crispmax) and $\epsilon$-orthogonality at convergence. \\

\noindent {\bf Lightweight connectivity with stochasticity.} Unless explicitly mentioned, $\A_k$ is simply rewritten as $\A$. We consider  a reparametrization   $\A = h(\hat{\A}) \D( h(\hat{\A}^\top))^{-1}$, with $\D(.)$ being the degree matrix operator, $h$ a strictly monotonic positive function and this allows a free setting of the matrix $\hat{\A}$ during optimization while guaranteeing stochasticity. In practice, $h$ is set to $\exp$ and the original gradient is obtained, similarly to Eq.~\ref{eq00001111}, from layerwise gradient back propagation by multiplying the original gradient by the Jacobian  $[\JJ_{\textrm{stc}}]_{ij,i'j'}= [{\A}_{i'j'}.(\delta_{ii'} - \A_{ij})]$ with $\delta_{ii'}=1_{\{i=i'\}}$. Note that stochasticity, when combined with orthogonality, leightens connectivity by a factor $n$ compared to orthogonality whose factor does not exceed $K$; this combination is obtained by multiplying the underlying Jacobians, so the final gradient becomes  
                            \begin{equation}\label{eq0000}
\displaystyle \frac{\partial \J}{\partial \v(\{\hat{\A}_k\}_k)} = \displaystyle   \JJ_\textrm{stc}. \JJ_\textrm{orth}. \frac{\partial \J}{\partial \v(\{{\A}_k\}_k)},
                                                                                                                                            \end{equation} 
and this order of application is strict, as orthogonality sustains after stochasticity  while the converse is not necessarily guaranteed at the end of the optimization process.

\section{Experiments}
{\bf Database and settings.} We evaluate the performance of our GCN on the task of action recognition using the First-Person Hand Action (FPHA) dataset~\cite{garcia2018}. The latter includes 1175 skeletons belonging to 45 action categories which are performed by 6 different individuals in 3 scenarios. Action categories are highly variable with inter and intra subject variability including style, speed, scale and viewpoint. Each video (sequence of skeletons) is initially described with a handcrafted graph $\G = (\V,\E)$ where each node $v_j \in \V$ corresponds to the $j$-th hand-joint trajectory (denoted as $\{\hat{p}_j^t\}_t$)  and an edge $(v_j, v_i) \in  \E$ exists iff the $j$-th and the $i$-th trajectories are spatially connected. Each trajectory in $\G$ is processed using {\it temporal chunking}: first, the total duration of a  sequence is split into $M$ equally-sized temporal chunks ($M=4$ in practice), then the trajectory  coordinates  $\{\hat{p}_j^t\}_t$  are assigned to the $M$ chunks (depending on their time stamps) prior to concatenate the averages of these chunks; this produces the raw description of $v_j$, again denoted as $\psi(v_j)$. \\
\noindent {\bf Implementation details.} We trained the GCNs end-to-end using the Adam optimizer for 2,800 epochs  with a batch size equal to $600$, a momentum of $0.9$ and a global learning rate (denoted as $\nu(t)$)  inversely proportional to the speed of change of the cross entropy loss used to train our networks; when this speed increases (resp. decreases),   $\nu(t)$  decreases as $\nu(t) \leftarrow \nu(t-1) \times 0.99$ (resp. increases as $\nu(t) \leftarrow \nu(t-1) \slash 0.99$). In all these experiments, we use a GeForce GTX 1070 GPU device (with 8 GB memory), we evaluate the performances using the 1:1 setting proposed in~\cite{garcia2018} with 600 action sequences for training and 575 for testing, and we report the average accuracy over all the classes of actions.\\

\noindent {\bf Performances and comparison.} We compare the performances of our GCN  design against two baselines: handcrafted and learned. In the first baseline (known as power map), all the matrices  $\{ \A_k\}_{k}$  are evaluated upon the adjacency matrix $\A$ (taken from the input skeletons) as $\A_k = \A^{(k)}$ with $\A^{(k)} = \A^{(k-1)} \A$, $\A^{(0)} = \I$  and this defines nested supports for convolutions while  in the second baseline, all the adjacency matrices $\{ \A_k\}_{k}$ are learned using the objective function~(\ref{matrixform3}) but w/o orthogonality and stochasticity constraints. Table~\ref{table21} shows a comparison with these baselines and an ablation study of our complete model and the impact of orthogonality (separately and combined) on the performances. These results show that orthogonality has a clear and a consistent positive impact on the performances  while stochasticity (when combined with orthogonality) provides lightweight GCNs with an extra gain in accuracy. Clearly, these two constraints act as regularizers that also reduce the number of training parameters thereby leading to highly effective and also efficient GCNs. In order to further investigate the impact of these two constraints, we compare the underlying GCNs against lightweight ones obtained differently, with magnitude pruning; the latter consists first in zeroing the smallest parameters in the learned GCNs, and then fine-tuning the remaining parameters. As shown in table~\ref{table21},  lightweight GCNs, trained with orthogonality and stochasticity, clearly outperform those obtained with magnitude pruning+fine-tuning. Finally, we compare the classification performances of our GCN against other related methods in action recognition ranging from sequence based such as LSTM to deep graph (non-vectorial) methods, etc. (see table~\ref{compare2} and references within). From the results in these tables, our GCN brings a noticeable gain w.r.t. related state of the art methods.

 \begin{table}[ht]
 \begin{center}
\resizebox{0.69\columnwidth}{!}{
\begin{tabular}{cc||c|c||c|c||c|c}
  &  & \rotatebox{55}{H} & \rotatebox{55}{L} &     \rotatebox{55}{L+orth} & \rotatebox{55}{L+MP} &  \rotatebox{55}{L+orth+stc} & \rotatebox{55}{L+MP} \\
 \hline
  \hline
  \multirow{2}{*}{\rotatebox{35}{$K=2$}}& Accuracy (\%) &    84.17 & 83.30  &  84.52     & 84.52   & 83.65    & 81.56 \\
 &   Pruning rate (\%) &   none       &   none         &     50        &  50    &   95        & 95\\
  \hline 
   \multirow{2}{*}{\rotatebox{35}{$K=4$}} & Accuracy (\%) &   82.95 & 83.82  &  85.21 &  83.13  & 85.73 & 82.95\\
  & Pruning rate (\%) &   none        &   none         & 75         & 75     &  95           & 95  \\
  \hline 
  \multirow{2}{*}{\rotatebox{35}{$K=8$}}& Accuracy (\%) &  72.69 &  83.82  &  85.04  & 83.65 & \bf86.78 & 84.00 \\
  & Pruning rate (\%) &    none        &   none         &   87          &  87    &    95        & 95\\
  \end{tabular}}
\end{center}
\caption{Detailed performances, for different $K$, using handcrafted and learned connectivity w/o and with our constraints. We also compare these results with those of GCNs obtained using magnitude pruning (for the same pruning rates: $\floor{(1-\frac{1}{K}) \times 100}$ for L+orth vs. L+MP and $\floor{(1-\frac{1}{n})\times 100}$ for L+orth+stc vs. L+MP), here  H, L, orth, stc and MP stands respectively for handcrafted, learned, orthogonality, stochasticity and magnitude pruning.}\label{table21}
\end{table}

\begin{table}[ht]
\begin{center}
  \resizebox{0.59\columnwidth}{!}{
\begin{tabular}{cccc|c}
{\bf Method} & {\bf Color} & {\bf Depth} & {\bf Pose} & {\bf Accuracy (\%)}\\
\hline
  Two stream-color \cite{refref10}   & \cmark  &  \xmark  & \xmark  &  61.56 \\
Two stream-flow \cite{refref10}     & \cmark  &  \xmark  & \xmark  &  69.91 \\ 
Two stream-all \cite{refref10}      & \cmark  & \xmark   & \xmark  &  75.30 \\
\hline 
HOG2-depth \cite{refref39}        & \xmark  & \cmark   & \xmark  &  59.83 \\    
HOG2-depth+pose \cite{refref39}   & \xmark  & \cmark   & \cmark  &  66.78 \\ 
HON4D \cite{refref40}               & \xmark  & \cmark   & \xmark  &  70.61 \\ 
Novel View \cite{refref41}          & \xmark  & \cmark   & \xmark  &  69.21  \\ 
\hline
1-layer LSTM \cite{Zhua2016}        & \xmark  & \xmark   & \cmark  &  78.73 \\
2-layer LSTM \cite{Zhua2016}        & \xmark  & \xmark   & \cmark  &  80.14 \\ 
\hline 
Moving Pose \cite{refref59}         & \xmark  & \xmark   & \cmark  &  56.34 \\ 
Lie Group \cite{Vemulapalli2014}    & \xmark  & \xmark   & \cmark  &  82.69 \\ 
HBRNN \cite{Du2015}                & \xmark  & \xmark   & \cmark  &  77.40 \\ 
Gram Matrix \cite{refref61}         & \xmark  & \xmark   & \cmark  &  85.39 \\ 
TF    \cite{refref11}               & \xmark  & \xmark   & \cmark  &  80.69 \\  
\hline 
JOULE-color \cite{refref18}         & \cmark  & \xmark   & \xmark  &  66.78 \\ 
JOULE-depth \cite{refref18}         & \xmark  & \cmark   & \xmark  &  60.17 \\ 
JOULE-pose \cite{refref18}         & \xmark  & \xmark   & \cmark  &  74.60 \\ 
JOULE-all \cite{refref18}           & \cmark  & \cmark   & \cmark  &  78.78 \\ 
\hline 
Huang et al. \cite{Huangcc2017}     & \xmark  & \xmark   & \cmark  &  84.35 \\ 
Huang et al. \cite{ref23}           & \xmark  & \xmark   & \cmark  &  77.57 \\  

  \hline
Our best (table~\ref{table21})                    & \xmark  & \xmark   & \cmark  & \bf86.78                                                 
\end{tabular}}
\caption{Comparison against  state of the art methods.}\label{compare2}
\end{center}
 \end{table} 

\section{Conclusion} 
We introduce in this paper a novel framework that designs graph topology as a part of an ``end-to-end'' GCN learning. This topology is captured using multiple adjacency matrices whose optimization is constrained with orthogonality and stochasticity. The former makes it possible to remove the redundancy while the latter allows learning lightweight and   highly effective GCNs. These two constraints also act as regularizers that model structural relationships between network parameters in order to enhance both their  generalization and lightweightness. Experiments conducted on the challenging task of skeleton-based hand-gesture   recognition, shows the outperformance of the proposed lightweight GCNs against different baselines as well as the related work.

\end{document}